Chair: Pradyumna

Next meeting chair: Lim

Attended: Jasper, Simin, Lim, Jay, Jamie, David, David Smith

Apologies: Eloise

General Issues:

  • Check proc file to recognize the necessary samplers for the experiment.
  • The Y-pol issue at Ke bit of a problem with aua/aov. Air conditioner issues too. Hard drive fail on flexbuffke. Currently inoperable. Jamie to send an update.

Handover issues:

  • AUA085: Yg - mk5 failed to recognize disk name. On-site reboot fixed it; Ke - Y polarization is cactus? -issues with Ypol; Hb - fs crashed which required a hard restart - probably a power glitch/bad UPS.
  • AOV068: Hb - MET sensor issue? -diagnosed and fixed. wth should return temperature/pressure readings; Ke - Stats for B/D (Y-pol) intermittently bad from start of session - probably due to broken/cracked cable. Unable to fix remotely. Removed disk6 (I/O) errors. set_disks? now reports 35 drives.